OpenStudy (anonymous):

A tent is shaped like a pyramid. It has a square base measuring 8ft on a side and a height of 6ft. What is the volume of the tent?

OpenStudy (anonymous):

I got 128 ft^3 I dont know if its right or not

OpenStudy (anonymous):

the equation for the area is 1/3(area of base)(height) so for this problem it would be: 1/3 (8*8)(6) = 1/3(64)(6) = 192 ft^3

OpenStudy (anonymous):

Thats not one of the aswers to choose from. They are 512ft^3 128ft^3 16ft^3 and 32ft^3

OpenStudy (anonymous):

my bad the answer is 128ft^3. i accidentally divided the 6 by two instead of 3. you're right.

OpenStudy (anonymous):

ok thanks